OpenStudy (anonymous):

What are the coordinates of the turning point for the function f(x) = (x + 2)3 - 4?

OpenStudy (anonymous):

is answer 1 ?

OpenStudy (anonymous):

-2/3 and 1.because when x=-2/3 > y=0 and x=1 > y is positive

OpenStudy (anonymous):

but i am not sure it is answer that you looking for

OpenStudy (anonymous):

(-2, -4) (-2, 4) (2, -4) (2, 4)

OpenStudy (anonymous):

the turning point means vertex right?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

is the equation actually written like this? f(x) = (x + 2)^3 - 4

OpenStudy (precal):

turning point is where it changes concavity from concave up to concave down

OpenStudy (precal):



OpenStudy (precal):

\[y=x^3\]

OpenStudy (anonymous):

yes

OpenStudy (precal):

turning point is origin because it goes from concave down to concave up

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

if so it would be -2,-4

OpenStudy (precal):

is this calculus?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

more like algrbra 2

OpenStudy (anonymous):

thank you

OpenStudy (precal):

ok (-2,-4) you could use transformations

OpenStudy (anonymous):

What are the zeros of the polynomial function f(x) = x3 - 2x2 - 24x?

OpenStudy (precal):

parent function is at (0,0) and it got moved 2 units left and down 4

OpenStudy (precal):

now you have to solve for where the function crosses the x axis. It is to the third power and you have 3 solutions maximum. Factor out your GCF and then factor again. You will see it, or look at the graph and determine at what values it crosses the x axis

OpenStudy (anonymous):

-6, 0, 4 6, 0, -4 -6, 0, -4 6, 0, 4

OpenStudy (precal):

sorry, seems like you just want the solutions and are not interested in the process. Learning the process will help you not only today but as you progress in more advance classes.

OpenStudy (anonymous):

autum all you need to do is factor the polynomial and set it to equal 0, you will find your answer that way.

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

did you factor and set it to equal 0?

OpenStudy (anonymous):

yes did i get it right?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

what did you get when you factored?

OpenStudy (anonymous):

6,0,-4

OpenStudy (anonymous):

you have to factor the funtion, those are the zeros